The markup on a television set is 20 percent of the cost. The markup is what percent of the selling price?
(markup = selling price - cost)
(A) 8%
(B) 10%
(C) 12.5%
(D) 15%
(E) 16.66%


Could you please kindly provide me the solution of the above problem.

Thanks

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

The markup on a television set is 20 percent of the cost. The markup is what percent of the selling price?
(markup = selling price - cost)
---
Let cost = "x" dollars.
Then markup is "x+0.20x" = 1.2x
-----
To find percent the markup is to the selling price
0.20x/1.2x = 0.20/1.2 = 0.1667 = 16.67%
--------------
Cheers,
Stan H.
=================
(A) 8%
(B) 10%
(C) 12.5%
(D) 15%
(E) 16.66%

Answer by solver91311(24713)



Let represent the cost. Let represent the selling price. We are given that the markup is 20% of the cost, so the markup is . From that we can write:



And a little algebra...



The markup as a percentage of sell price would be given by markup divided by sell price, so:



You can do your own arithmetic.


John

My calculator said it, I believe it, that settles it